OpenStudy (anonymous):

find the solutions for the equation below. Check all that apply. 2x^2+9x+4=0 ax=-3 bx=7 c=-1/2 d.x=-1/2 e.x=2 f.x=4

OpenStudy (anonymous):

\[(x+4)(2x+1)=0 \]

OpenStudy (anonymous):

got it its c and 4

OpenStudy (anonymous):

c ND F

OpenStudy (anonymous):

Yes? Is F = -4 ?

OpenStudy (anonymous):

nope

OpenStudy (radar):

Then forget about f.

OpenStudy (anonymous):

its suppose to be d.x=4

OpenStudy (anonymous):

i mean d x=-4

OpenStudy (anonymous):

wait so is it c and f or c and d

OpenStudy (radar):

What does d. read as?

OpenStudy (radar):

d. is a better answer than c.

OpenStudy (radar):

c. doesn't say what is equal to -1/2

OpenStudy (anonymous):

But it says D = -4?

OpenStudy (anonymous):

@romanortiz65 CAN YOU WRITE THE ANSWER CHOICES AGAIN?

OpenStudy (radar):

If d. reads as " x = -4" then that is good also.

OpenStudy (anonymous):

ax=-3 bx=7 c=-1/2 d.x=-4 e.x=2 f.x=4

OpenStudy (anonymous):

Oh gosh. I'm looking at the wrong problem... Yes its is C and D

OpenStudy (anonymous):

im going offline now :P

OpenStudy (radar):

Select which of these are correct: A. bx=-36 B. cx=-2 C. x= -2.

OpenStudy (radar):

Never mind. Good luck, that was for @romanortiz65